Two iceboats A and B race on a frictionless lake. A has a mass and of m and B has mass of 2m. The wind exerts a constant force F on both boats along their direction of motion. Both boats start from rest , and both travel at the same distance s to finish line.At the finish line the speed of A is (a)equal to the speed of B
(b)greater than the speed of B
(c)less than the speed design of B
(d) cannot be obtained from the information given
Explanation of answer please
1
Expert's answer
2019-10-17T11:31:27-0400

For A:


"a_A=\frac{F}{m}"

"v_A^2-0^2=2a_As=\frac{2sF}{m}"

"v_A=\sqrt{\frac{2sF}{m}}"


For B:


"a_B=\frac{F}{2m}"

"v_B^2-0^2=2a_Bs=\frac{sF}{m}"

"v_B=\sqrt{\frac{sF}{m}}"

Answer: (b)greater than the speed of B